# Groups Acting Acylindrically on Trees

### v5

@@ 1. Introduction
The definition of an acylindrical action on a tree was first formulated by Sela [Sela97], and later generalised by Weidmann [Weidmann07] to the following.
@@
A third definition was formulated by Bowditch [Bowditch08] to give a property of an action on a general metric space, which is more coarse-geometric. The definition of Bowditch can be shown to agree with that of Weidmann in the context of actions on trees, in a result essentially due to of Minasyan and Osin [minasyanOsin13](c.f. Theorem 2.17). Groups that act acylindrically non-elementarily (c.f. Theorem 2.3) on hyperbolic spaces are a well-studied generalisation of hyperbolic groups. Such groups are called acylindrically hyperbolic.
The class of acylindrically hyperbolic groups is very broad – such a large number and variety of groups are acylindrically hyperbolic that it can be difficult to discern whether they have certain properties in general. For example, it is still an open problem whether the property of being acylindrically hyperbolic is preserved under quasi-isometry [Osin17, Question 2.20(a)] or even passes to finite index overgroups [minasyanOsin19, Question 2], although some progress was made on the latter question by Balasubramanya [BalasubramanyaFiniteIndex].
It is natural to ask which acylindrically hyperbolic groups act acylindrically non-elementarily on simplicial trees. Indeed, it is a well known result of Balasubramanya that all acylindrically hyperbolic groups admit acylindrical actions on quasi-trees [Bala16], and further restricting to trees allows us to make use of a wide range of strong theory. All actions will be by isometry, and all trees will be assumed to be simplicial with the natural edge metric. We make the following definition.
@@
Many groups are already known to be acylindrically arboreal, although we believe that this terminology is novel. For example Wilton and Zalesskii show that if a closed and orientable irreducible 3-manifold $M$ is non-geometric then the splitting of $\pi_{1}(M)$ given by the JSJ-decomposition of $M$ induces an acylindrical action on the associated Bass–Serre tree [WiltonZalesskii08, Lemma 2.4], so such groups will either be acylindrically arboreal or virtually cyclic by [Osin13, Theorem 1.1] (c.f. Theorem 2.3).
@@
The vast majority of other examples arise from a result of Gitik, Mj, Rips and Sageev [GitikMahanRipsSageev98, Main Theorem] that implies that a splitting of a hyperbolic group over a quasi-convex subgroup will induce an acylindrical action on the associated Bass–Serre tree. The study of quasi-convex splittings of hyperbolic groups has proved very fruitful (see [HaglundWise12, WiseQCBook], for example).
@@
The acylindrical action on a tree rather than on a generic hyperbolic metric space makes studying the properties of acylindrically arboreal groups much simpler than studying the properties of acylindrically hyperbolic groups in general. For example, the following are properties of acylindrically arboreal groups that are either unknown for acylindrically hyperbolic groups in general (even for those which split) or have a much more satisfying solutions in our restricted class.
@@
  • Quasi-isometry : It is very simple to see that acylindrical arboreality is not preserved by quasi-isometry. For example, the $(2,3,7)$-triangle group has property (FA) (c.f. Definition 2.13) by a well known result of Serre ([stilwell2002trees, Corollary 2 of Theorem I.6.26], c.f. Theorem 2.11), and this property will naturally prohibit any action on a tree from fulfilling the non-elementary requirement in the definition of acylindrical arboreality. However, this group is virtually a hyperbolic surface group, which is acylindrically arboreal by considering cuts along simple closed curves.
@@
  • Growth Rates : Using the fact that in a non-elementary acylindrical action on a tree we are guaranteed to find loxodromic elements in words of length at most two over any generating set (c.f Corollary 2.12), we can learn much about the growth properties of acylindrically arboreal groups. In particular, Kerr [Kerr2021ProductSG, Proposition 1.0.9] and Fujiwara [Fujiwara21, Theorem 1.1] prove strong results about the growth rates of acylindrically hyperbolic groups where we can find loxodromics quickly. It is however worth noting that the Fujiwara result requires equational Noetheriality in an essential way, and it is still very much open when an acylindrically arboreal group is equationally Noetherian (see [Valiunas20, Theorem 1.9] for a strong condition under which equational Noetheriallity will hold).
@@
  • Explicit Constructions in Bounded Cohomology : Monod and Shalom explicitly use the acylindrical action on a tree to show that all acylindrically arboreal groups are in the class $\mathcal{C}_{reg}$ ([MonodShalom04, Theorem 7.7, Corollary 7.10 and Remark 7.11]), which is equivalent to the second bounded cohomology with coefficients in the regular representation not vanishing (see [MonodShalom06] for a formal treatment of this property). Their proof is via an explicit construction of the desired non-boundary cocycle, and although subsequent results have shown that all acylindrically hyperbolic groups are in $\mathcal{C}_{reg}$ ([Hamenstadt08, Corollary B], [HullOsin13, Corollary 1.7]), the constructions in the general case are much less explicit.
@@
An acylindrical action on a tree also has strong implications for the Farrel–Jones conjecture, as shown by Knopf [Knopf19].
@@
Our main results are complete classifications of two very important classes of groups, those of graph products of groups and fundamental groups of compact and orientable hyperbolic 3-manifolds with empty or toroidal boundary. The acylindrical hyperbolicity of groups in these classes was considered in [minasyanOsin13], and while all manifolds considered in this paper have relatively hyperbolic fundamental groups which must then be acylindrically hyperbolic [Osin13, Theorem 1.2 and Proposition 2.12], the results of this paper relating to graph products of groups should be compared to those in [minasyanOsin13, Section 2.3].
@@
In a similar vein, we prove a similar sufficient condition in Corollary 3.13 for a subgroup of graph product of groups to be acylindrical arboreality, which should be compared with [minasyanOsin13, Theorem 2.12].
@@
This is similar to the definition of property $\mathbb{R}$ defined by Culler and Vogtmann [Culler1996AGT] for groups acting on real trees, excepting that they only allow the action of the group on the fixed line to be by translation.
We will see that property (FA-) prohibits acylindrical arboreality in the sense that a group with (FA-) can never fulfil the non-elementary condition when acting upon a tree, but that the reverse implication is not true — there exist many acylindrically hyperbolic (or even hyperbolic) groups that do not have (FA-) but which are not acylindrically arboreal. For example, let $Q$ be a finitely presented but not hyperbolic group with interesting actions on trees, such as the Baumslag Solitar group $BS(2,3)\cong\langle a,b\mid ab^{2}a^{-1}b^{-3}\rangle$. Using the Rips construction of Ollivier and Wise [OlivierWise04, Theorem 1.1] it is possible to extend $Q$ by an infinite group $N$ with Kazhdan’s property (T) (see [Zuk03], for example) to obtain a hyperbolic group $G$, which will have an infinite normal (T) (and therefore (FA-) by [Watatani], or [NibloReeves97, Theorem B]) subgroup. Such a subgroup prohibits the acylindrical arboreality of $G$ by the fact that as with acylindrical hyperbolicity, acylindrical arboreality is inherited by infinite normal subgroups (Lemma 2.18(3), see [Osin13, Lemma 7.1] for the original proof for acylindrically hyperbolic groups). However, $G$ will not (FA)- due to the projection to $Q$, which has an action on a tree with no fixed points or lines arising from the natural HNN structure.
@@
Our second main theorem provides a complete classification of when the fundamental group of a compact and orientable hyperbolic 3-manifold with empty or toroidal boundary is acylindrically arboreal.
@@
Along with the proof of a result of Minasyan and Osin [minasyanOsin19, Theorem 3], this theorem may be used to give a complete classification of the acylindrical arboreality of the fundamental groups of closed and orientable $3$-manifolds.
@@
A negative answer to Question 1.7 would require a rather pathological construction, although we are inclined to believe that the answer will be in the negative. For example, the double of a hyperbolic group $G$ over a finitely presented subgroup $H$ will be hyperbolic only if $H$ is undistorted [BridsonHaefliger, Theorem I.$\Gamma$6.20], so a counterexample using the doubling construction would require a subgroup of a hyperbolic group which is finitely generated but not finitely presented. Such examples are poorly understood in general, and tend to have large normalisers in the ambient group [WiseCoherence].
@@ 2.1. Acylindrical Hyperbolicity
We will define a hyperbolic metric space to be a geodesic metric space satisfying any of the standard equivalent conditions, for example the thin triangles condition for $\delta$-hyperbolicity due to Rips [BridsonHaefliger, Definition III.H.1.1]. As in the introduction we have the following coarse-geometric definition of acylindricity due to Bowditch.
@@
Let $G$ be a group acting on a metric space $(X,d)$ by isometry. For all $\epsilon\geq 0$, $x,y\in X$ we define the Pointwise Quasi-Stabiliser $\operatorname{PQStab}_{G}^{\epsilon}(\{x,y\})$ to be the set
$\{g\in G\mid d(x,gx)\leq\epsilon\text{ and }d(y,gy)\leq\epsilon\}.$
We say that $G$ acts acylindrically on $X$ if for all $\epsilon\geq 0$ there exist constants $R(\epsilon)$, $N(\epsilon)\geq 0$ such that for all $x,y\in X$ with $d(x,y)\geq R(\epsilon)$, we have $|\operatorname{PQStab}_{G}^{\epsilon}|\leq N(\epsilon)$.
@@
Let $G$ be a group acting on a hyperbolic space $X$ by isometry, and let $g\in G$. Then we say that $g$:
  • is elliptic if $g$ has bounded orbits;
  • is loxodromic if for any $x\in X$ the map $\mathbb{Z}\rightarrow X$ given by $n\mapsto g^{n}x$ is a quasi-isometric embedding;
We will require some results pertaining to acylindrically hyperbolic groups, which we will include without proof for brevity. We first have the following classification theorem due to Osin. A similar classification for generic actions on hyperbolic spaces was originally proved by Gromov in [Gromov87, Section 8], with more possibilities, but these may be disregarded as they will not occur in the acylindrical case.
@@
[Osin13, Theorem 1.1] Let $G$ be a group acting acylindrically on a hyperbolic space $X$. Then $G$ satisfies exactly one of the following conditions.
@@
Let $G$ be be a group. We say that $G$ is acylindrically hyperbolic if $G$ admits a non-elementary acylindrical action on a hyperbolic space $X$.
@@
Further examples include the mapping class groups MCG$(\Sigma_{g,p})$ of closed surfaces of genus $g$ with $p$ punctures unless $g=0$ and $p\leq 3$, $Out(F_{n})$ for $n\geq 2$ and one relator groups with at least three generators (see [Osin13, Section 8], for example).
@@ 2.2. Groups Acting on Trees
Let $\Gamma=(V(\Gamma),E(\Gamma))$ be a graph. We say that $\Gamma$ is finite if $|V(\Gamma)|<\infty$, and we say that $\Gamma$ is simple if $E(\Gamma)$ contains no loops or multiedges.
@@
    We define the $n$-path $P_{n}$ for $n\geq 2$ to be the unique (up to isomorphism) connected graph on $n$ vertices with $n-1$ edges and maximum vertex degree two, and the $n$-cycle $C_{n}$ for $n\geq 3$ to be the unique (up to isomorphism) connected graph on $n$ vertices with $n$ edges such that the degree of every vertex is two.
@@
We will assume the reader has some familiarity with Bass–Serre theory, and for a more detailed discussion we refer to [stilwell2002trees, dicks2011groups]. Let $(\Gamma,\mathfrak{G})$ be a graph of groups, where $\Gamma$ is a connected directed graph that may not be finite or simple and $\mathfrak{G}$ is the following data:
@@
  • To every edge $e\in E(\Gamma)$ we assign monomorphisms $d_{0}:G_{e}\rightarrow G_{i(e)}$ and $d_{1}:G_{e}\rightarrow G_{t(e)}$, where $i(e)$ and $t(e)$ are the initial and terminal vertices of $e$ in $\Gamma$ respectively.
We will use a slight abuse of notation to consider each vertex group $G_{v}$ as a subgroup the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ along the natural inclusion. Similarly, we will consider each edge group $G_{e}$ to be the subgroup of the fundamental group given by the image of $d_{0}(G_{e})$ in the vertex group $G_{i(e)}$. We call a graph of groups is trivial if there exists some $v\in V(\Gamma)$ such that $G_{v}=\pi_{1}(\Gamma,\mathfrak{G})$, or non-trivial otherwise. We say that a graph of groups $(\Gamma,\mathfrak{G})$ is a graph of groups decomposition or splitting of a group $G$ if the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ is isomorphic to $G$. We denote by $T(\Gamma,\mathfrak{G})$ the Bass–Serre tree associated to the splitting, on which $G$ acts naturally by isometry with respect to the edge metric and without inversion [stilwell2002trees, Section I.5.3].
@@
The following lemma is a restating of Proposition I.2.10 of [stilwell2002trees] and its corollary which allows us to remove any ambiguity as to the definition of an elliptic action in the context of simplicial trees and which we include without proof.
@@
It follows from the fundamental theorem of Bass–Serre theory that an action is elliptic if and only if the quotient graph of groups is trivial.
@@
[stilwell2002trees, Corollary 2 of Theorem I.6.26] Let $G$ be a group generated by a finite number of elements $s_{1},...,s_{m}$ and let $T$ be a tree on which $G$ acts by isometry. Assume further that for all for distinct all pairs $1\leq i,j\leq m$ we have that $s_{i},s_{j}$ and $s_{i}s_{j}$ act elliptically on $T$. Then $G$ must act elliptically on $T$.
@@
Note that this will hold even in the case when $S$ is uncountable, as any loxodromic element would have to appear as finite word over the elements of $S$, so would sit in a finitely generated subgroup where we can apply Theorem 2.11.
@@
We have the following definition due Serre [stilwell2002trees, Section I.6.1].

### v6

@@ 1. Introduction
The definition of an acylindrical action on a tree was first formulated by Sela [undefaj], and later generalised by Weidmann [undefao] to the following.
@@
A third definition was formulated by Bowditch [undefd] to give a property of an action on a general metric space, which is more coarse-geometric. The definition of Bowditch can be shown to agree with that of Weidmann in the context of actions on trees, in a result essentially due to Minasyan and Osin [undefz](c.f. Theorem 2.17). Groups that act acylindrically non-elementarily (c.f. Theorem 2.3) on hyperbolic spaces are a well-studied generalisation of hyperbolic groups. Such groups are called acylindrically hyperbolic.
The class of acylindrically hyperbolic groups is very broad – such a large number and variety of groups are acylindrically hyperbolic that it can be difficult to discern whether they have certain properties in general. For example, it is still an open problem whether the property of being acylindrically hyperbolic is preserved under quasi-isometry [undefag, Question 2.20(a)] or even passes to finite index overgroups [undefaa, Question 2], although some progress was made on the latter question by Balasubramanya [undefc].
It is natural to ask which acylindrically hyperbolic groups act acylindrically non-elementarily on simplicial trees. Indeed, it is a well-known result of Balasubramanya that all acylindrically hyperbolic groups admit acylindrical actions on quasi-trees [undefb], and further restricting to trees allows us to make use of a wide range of strong theory. All actions will be by isometry, and all trees will be assumed to be simplicial with the natural edge metric. We make the following definition.
@@
Many groups are already known to be acylindrically arboreal, although we believe that this terminology is novel. For example Wilton and Zalesskii show that if a closed and orientable irreducible 3-manifold $M$ is non-geometric then the splitting of $\pi_{1}(M)$ given by the JSJ-decomposition of $M$ induces an acylindrical action on the associated Bass–Serre tree [undefap, Lemma 2.4], so such groups will either be acylindrically arboreal or virtually cyclic by [undefaf, Theorem 1.1] (c.f. Theorem 2.3).
@@
The vast majority of other examples arise from a result of Gitik, Mj, Rips and Sageev [undefl, Main Theorem] that implies that a splitting of a hyperbolic group over a quasi-convex subgroup will induce an acylindrical action on the associated Bass–Serre tree. The study of quasi-convex splittings of hyperbolic groups has proved very fruitful (see [undefo, undefaq], for example).
@@
The acylindrical action on a tree rather than on a generic hyperbolic metric space makes studying the properties of acylindrically arboreal groups much simpler than studying the properties of acylindrically hyperbolic groups in general. For example, the following are properties of acylindrically arboreal groups that are either unknown for acylindrically hyperbolic groups in general (even for those which split) or have a much more satisfying solution in our restricted class.
@@
  • Quasi-isometry : It is very simple to see that acylindrical arboreality is not preserved by quasi-isometry. For example, the $(2,3,7)$-triangle group has property (FA) (c.f. Definition 2.13) by a well-known result of Serre ([undefak, Corollary 2 of Theorem I.6.26], c.f. Theorem 2.11), and this property will naturally prohibit any action on a tree from fulfilling the non-elementary requirement in the definition of acylindrical arboreality. However, this group is virtually a hyperbolic surface group, which is acylindrically arboreal by considering cuts along simple closed curves.
@@
  • Growth Rates : Using the fact that in a non-elementary acylindrical action on a tree we are guaranteed to find loxodromic elements in words of length at most two over any generating set (c.f Corollary 2.12), we can learn much about the growth properties of acylindrically arboreal groups. In particular, Kerr [undefv, Proposition 1.0.9] and Fujiwara [undefk, Theorem 1.1] prove strong results about the growth rates of acylindrically hyperbolic groups where we can find loxodromics quickly. It is however worth noting that the Fujiwara result requires equational Noetheriality in an essential way, and it is still very much open when an acylindrically arboreal group is equationally Noetherian (see [undefam, Theorem 1.9] for a strong condition under which equational Noetheriality will hold).
@@
  • Explicit Constructions in Bounded Cohomology : Monod and Shalom explicitly use the acylindrical action on a tree to show that all acylindrically arboreal groups are in the class $\mathcal{C}_{\operatorname{reg}}$ (see [undefab, Theorem 7.7, Corollary 7.10 and Remark 7.11]), which is equivalent to the second bounded cohomology with coefficients in the regular representation not vanishing (see [undefac] for a formal treatment of this property). Their proof is via an explicit construction of the desired non-boundary cocycle, and although subsequent results have shown that all acylindrically hyperbolic groups are in $\mathcal{C}_{\operatorname{reg}}$ (see [undefp, Corollary B], [undefu, Corollary 1.7]), the constructions in the general case are much less explicit.
@@
An acylindrical action on a tree also has strong implications for the Farrel–Jones conjecture, as shown by Knopf [undefx].
@@
Our main results are complete classifications of two very important classes of groups, those of graph products of groups and fundamental groups of compact and orientable hyperbolic 3-manifolds with empty or toroidal boundary. The acylindrical hyperbolicity of groups in these classes was considered in [undefz], and while all manifolds considered in this paper have relatively hyperbolic fundamental groups which must then be acylindrically hyperbolic [undefaf, Theorem 1.2 and Proposition 2.12], the results of this paper relating to graph products of groups should be compared to those in [undefz, Section 2.3].
@@
In a similar vein, we prove a similar sufficient condition in Corollary 3.13 for a subgroup of a graph product of groups to be acylindrical arboreal, which should be compared with [undefz, Theorem 2.12].
@@
[…]
We will see that property (FA-) prohibits acylindrical arboreality in the sense that a group with (FA-) can never fulfil the non-elementary condition when acting upon a tree, but that the reverse implication is not true — there exist many acylindrically hyperbolic (or even hyperbolic) groups that do not have (FA-) but which are not acylindrically arboreal. For example, let $Q$ be a finitely presented but not hyperbolic group with an action on a tree with no fixed points or lines, such as the Baumslag–Solitar group $\operatorname{BS}(2,3)\cong\langle a,b\mid ab^{2}a^{-1}b^{-3}\rangle$. Using the Rips construction of Ollivier and Wise [undefae, Theorem 1.1] it is possible to extend $Q$ by an infinite group $N$ with Kazhdan’s property (T) (see [undefas], for example) to obtain a hyperbolic group $G$, which will have an infinite normal (T) (and therefore (FA-) by [undefan], or [undefad, Theorem B]) subgroup. Such a subgroup prohibits the acylindrical arboreality of $G$ by the fact that as with acylindrical hyperbolicity, acylindrical arboreality is inherited by infinite normal subgroups (Lemma 2.18(3), see [undefaf, Lemma 7.1] for the original proof for acylindrically hyperbolic groups). However, $G/N\cong Q$ admits an action on a tree with no fixed points or lines by construction, so $G$ will not have (FA)- as it will act on the same tree via the quotient map.
@@
Our second main theorem, Theorem 1.6 below, provides a complete classification of when the fundamental group of a compact and orientable hyperbolic 3-manifold with empty or toroidal boundary is acylindrically arboreal — see Section 4 for the relevant definitions.
@@
Along with the proof of a result of Minasyan and Osin [undefaa, Theorem 3], Theorem 1.6 restricted to the closed case may be used to give a complete classification of the acylindrical arboreality of the fundamental groups of closed and orientable $3$-manifolds.
@@
[…]
@@ 2.1. Acylindrical Hyperbolicity
We will define a hyperbolic metric space to be a geodesic metric space satisfying any of the standard equivalent conditions, for example the thin triangles condition for $\delta$-hyperbolicity due to Rips [undefe, Definition III.H.1.1]. We have the following coarse-geometric definition of acylindricity due to Bowditch.
@@
Let $G$ be a group acting on a metric space $(X,d)$ by isometry. For all $\epsilon\geq 0$, $x,y\in X$ we define the Pointwise Quasi-Stabiliser $\operatorname{PQStab}_{G}^{\epsilon}(\{x,y\})$ to be the set $\{g\in G\mid d(x,gx)\leq\epsilon\text{ and }d(y,gy)\leq\epsilon\}.$ We say that $G$ acts acylindrically on $X$ if for all $\epsilon\geq 0$ there exist constants $R(\epsilon)$, $N(\epsilon)\geq 0$ such that for all $x,y\in X$ with $d(x,y)\geq R(\epsilon)$, we have $|\operatorname{PQStab}_{G}^{\epsilon}|\leq N(\epsilon)$.
@@
Let $G$ be a group acting on a hyperbolic space $X$ by isometry, and let $g\in G$. Then we say that
  • $g$ is elliptic if $g$ has bounded orbits;
  • $g$ is loxodromic if for any $x\in X$ the map $\mathbb{Z}\rightarrow X$ given by $n\mapsto g^{n}x$ is a quasi-isometric embedding;
We will require some results pertaining to acylindrically hyperbolic groups, which we will include without proof for brevity. We first have the following classification theorem due to Osin. A similar classification for generic actions on hyperbolic spaces was originally proved by Gromov in [undefn, Section 8], with more possibilities, but these may be disregarded as they will not occur in the acylindrical case.
@@
[undefaf, Theorem 1.1] Let $G$ be a group acting acylindrically on a hyperbolic space $X$. Then $G$ satisfies exactly one of the following conditions.
@@
Let $G$ be a group. We say that $G$ is acylindrically hyperbolic if $G$ admits a non-elementary acylindrical action on a hyperbolic space $X$.
@@
[…]
@@ 2.2. Groups Acting on Trees
Let $\Gamma=(V(\Gamma),E(\Gamma))$ be a graph. Then $\Gamma$ is said to befinite if $|V(\Gamma)|<\infty$, and we say that $\Gamma$ is simple if $E(\Gamma)$ contains no loops or multiedges.
@@
    We define the $n$-path $P_{n}$ for $n\geq 2$ to be the unique (up to isomorphism) connected graph on $n$ vertices with n1n-1 edges and maximum vertex degree two, and the $n$-cycle $C_{n}$ for $n\geq 3$ to be the unique (up to isomorphism) connected graph on $n$ vertices with $n$ edges such that the degree of every vertex is two.
@@
We will assume the reader has some familiarity with Bass–Serre theory, and for a more detailed discussion we refer to [undefak, undefi]. Let $(\Gamma,\mathfrak{G})$ be a graph of groups, where $\Gamma$ is a connected directed graph that may not be finite or simple and $\mathfrak{G}$ is the following data:
@@
  • To every edge $e\in E(\Gamma)$ we assign monomorphisms
    $d_{0}:G_{e}\rightarrow G_{i(e)}\text{ and }d_{1}:G_{e}\rightarrow G_{t(e)},$
    where $i(e)$ and $t(e)$ are the initial and terminal vertices of $e$ in $\Gamma$ respectively.
We will use a slight abuse of notation to consider each vertex group $G_{v}$ as a subgroup of the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ along the natural inclusion. Similarly, we will consider each edge group $G_{e}$ to be the subgroup of the fundamental group given by the image of $d_{0}(G_{e})$ in the vertex group $G_{i(e)}$. We call a graph of groups trivial if there exists some $v\in V(\Gamma)$ such that $G_{v}=\pi_{1}(\Gamma,\mathfrak{G})$, or non-trivial otherwise. We say that a graph of groups $(\Gamma,\mathfrak{G})$ is a graph of groups decomposition or splitting of a group $G$ if the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ is isomorphic to $G$. We denote by $T(\Gamma,\mathfrak{G})$ the Bass–Serre tree associated to the splitting, on which $G$ acts naturally by isometry with respect to the edge metric and without inversion [undefak, Section I.5.3].
@@
As in [undefak, Section I.5.4], an action on a tree will give rise to a quotient graph of groups decomposition $(T/G,\mathfrak{G}$) of $G$, where the vertex or edge group of a vertex or edge of $T/G$ is defined to have the isomorphism type of the stabiliser of any preimage of that vertex or edge in $T$, and the edge monomorphisms are defined similarly.
The following lemma is a restating of [undefak, Proposition I.2.10] and its corollary which allows us to remove any ambiguity as to the definition of an elliptic action in the context of simplicial trees and which we include without proof.
@@
It follows from the fundamental theorem of Bass–Serre theory that an action on a tree is elliptic if and only if the quotient graph of groups is trivial.
@@
[undefak, Corollary 2 of Theorem I.6.26] Let $G$ be a group generated by a finite number of elements $s_{1},...,s_{m}$ and let $T$ be a tree on which $G$ acts by isometry. Assume further that for all distinct pairs $1\leq i,j\leq m$ we have that $s_{i},s_{j}$ and $s_{i}s_{j}$ act elliptically on $T$. Then $G$ must act elliptically on $T$.
@@
Note that this will hold even in the case when $S$ is uncountable, as any loxodromic element would have to appear as a finite word over the elements of $S$, so would sit in a finitely generated subgroup where we can apply Theorem 2.11.
@@
We have the following definition due to Serre [undefak, Section I.6.1].